OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = –32 and a9 = –120?

OpenStudy (anonymous):

-373

OpenStudy (campbell_st):

you will need to find the common difference 1st the formula for a term in an arithmetic series is \[T _{n} = a + (n -1)d\] a= 1st term, n = number of terms and d = common differnece then \[-120 = -32 + (9 -1)d\] solve for d... then use the formula again with n = 32